# Reasoning in the Dark: Interleaved Vision-Text Reasoning in Latent Space

### v1

@@ Abstract
Multimodal reasoning aims to enhance the capabilities of MLLMs by incorporating intermediate reasoning steps before reaching the final answer. It has evolved from text-only reasoning to the integration of visual information, enabling the thought process to be conveyed through both images and text. Despite its effectiveness, current multimodal reasoning methods depend on explicit reasoning steps that require labor-intensive vision-text annotations and inherently introduce significant inference latency. To address these issues, we introduce multimodal latent reasoning with the advantages of multimodal representation, reduced annotation, and inference efficiency. To facilicate it, we propose Interleaved Vision-Text Latent Reasoning (IVT-LR), which injects both visual and textual information in the reasoning process within the latent space. Specifically, IVT-LR represents each reasoning step by combining two implicit parts: latent text (the hidden states from the previous step) and latent vision (a set of selected image embeddings). We further introduce a progressive multi-stage training strategy to enable MLLMs to perform the above multimodal latent reasoning steps. Experiments on M3CoT and ScienceQA demonstrate that our IVT-LR method achieves an average performance increase of 5.45% in accuracy, while simultaneously achieving a speed increase of over 5 times compared to existing approaches. Code available at https://github.com/FYYDCC/IVT-LR.
Reasoning in the Dark: Interleaved Vision-Text Reasoning in Latent Space
Chao Chen1, Zhixin Ma2, Yongqi Li1‡, Yupeng Hu3, Yinwei Wei3, Wenjie Li1, Liqiang Nie4 1The Hong Kong Polytechnic University 2Singapore Management University 3Shandong University 4Harbin Institute of Technology (Shenzhen) ochenchaoo@outlook.com liyongqi0@gmail.com
Corresponding author.
@@ 1 Introduction
Over the past few years, the capabilities of large language models (LLMs) have been further unlocked through advancements in reasoning. Researchers have sought to enhance the reasoning abilities of LLMs, initially through prompting techniques such as Chain-of-Thought prompting  Wei et al. (2022), and more recently by developing large reasoning models using reinforcement learning, like GPT-4o  Hurst et al. (2024) and Deepseek-R1  Guo et al. (2025). Building on the success of reasoning in LLMs, there has been growing interest in the research community to extend these reasoning capabilities to multimodal LLMs (MLLMs). This has brought the promising topic of multimodal reasoning, aiming to improve the performance of models on multimodal tasks, such as VQA, through reasoning.
@@
Current multimodal reasoning approaches can be broadly categorized into the following progressive steps: 1) Text-only reasoning. Early multimodal reasoning methods primarily focused on pure text-based reasoning, where MLLMs perform textual reasoning before generating the final answer. These approaches Hu et al. (2022); Mondal et al. (2024) could seamlessly apply LLMs methodologies to MLLMs. 2) Vision-text involved reasoning. Some studies  Hu et al. (2024); Liu et al. (2025); Chern et al. (2025) highlight that the intermediate reasoning steps also require the involvement of visual information. For instance, Gao et al. (2025) enhances reasoning by generating sequential steps that interleave visual information with textual rationales via selected image patches. In a related direction, Zheng et al. (2025) trains models through end-to-end reinforcement learning to autonomously zoom in on image regions for fine-grained visual inspection during the reasoning process. Alternatively, Li et al. (2025b) enables MLLMs to actively “think visually” by generating explicit image visualizations of their reasoning traces, thereby significantly enhancing performance on complex spatial reasoning tasks.
Recently, latent reasoning has emerged as a new paradigm in LLMs, which eliminates the need for explicit and lengthy textual reasoning by leveraging implicit latent vectors Hao et al. (2024). Inspired by this, we believe that latent reasoning holds even greater potential for facilitating vision-text interleaved intermediate reasoning steps due to the following reasons: 1) Multimodal representation potential. Latent reasoning enables the reasoning process to occur entirely within a hidden space, offering a greater capacity to represent rich, multimodal information during reasoning. 2) Reduced Annotation. Introducing latent reasoning will lessen reliance on heavily annotated vision-text interleaved reasoning data, as reasoning steps no longer need to be fully observable or linguistically aligned. 3) Inference efficiency. By avoiding long chains of explicit multimodal representation in the reasoning step, it will significantly improve efficiency.
@@
  • We introduce IVT-LR, the first framework to achieve fully multimodal latent reasoning. Unlike prior methods, our approach enables both textual and visual information to be reasoned with in the latent space, eliminating the need for intermediate explicit text or image generation.
@@ 2.1 Multimodal Reasoning
Text-only reasoning.
Early works attempt to convert visual information into text before reasoning, using tools or visual experts to generate textual representations to guide LLMs. Hu et al. (2022) first introduced the concept of captions, extracting visual content as textual captions and concatenating them to the input to enhance reasoning. Inspired by this, subsequent works pursued finer-grained understanding of images to improve textual expressiveness. Zheng et al. (2023) generates a rationale that incorporates image information from visual-text inputs, which is then used for reasoning. Other works Mitra et al. (2024); Mondal et al. (2024) leverage graph structures to identify entities in images and construct relationships among them, enhancing reasoning based on these inter-entity connections.
Vision-text involved reasoning.
This line of work emphasizes using images together with text during the rationale generation and reasoning process. Building on the reasoning paradigm of large language models, Zhang et al. (2024) first proposed decoupling rationale generation from answer generation in the Vision-Text Reasoning field. Subsequently, Shao et al. (2024) annotates key regions of the original image in intermediate steps, training models to focus on image regions relevant to the answer. While some works Gao et al. (2025); Zhang et al. (2025) further extract key image regions progressively during reasoning, combining visual information with textual reasoning to generate the final answer. Moreover, new methods Hu et al. (2024); Liu et al. (2025) emulate human thought by sketching images during reasoning, focusing on core concepts, structures, and relationships while ignoring redundant details. Other works Li et al. (2025b); Chern et al. (2025) generate new images during reasoning, combining them with text to improve reasoning in complex scenarios. To completely decouple reasoning from language and amplify the role of images, Xu et al. (2025) proposes reasoning solely with newly generated images, achieving substantial improvements in visual navigation tasks.
2.2 Latent Reasoning
Latent reasoning refers to internal, non-linguistic thinking performed in a hidden latent space before generating the final answer. Early methods used special tokens to guide latent reasoning. Goyal et al. (2024) introduces learnable <pause> tokens, giving the model opportunities to internally update information before generating an answer, while Wang et al. (2024b) uses <plan> tokens to guide reasoning.
Later, some works exploit the model’s continuous hidden states to replace explicit reasoning steps. Hao et al. (2024) pioneers continuous latent space reasoning by feeding the last hidden states as input embeddings for the next step without generating intermediate tokens, significantly reducing reasoning tokens and improving efficiency. Inspired by this, subsequent methods improve the quality of intermediate representations. Cheng and Van Durme (2024) uses variable-length contemplation tokens for latent reasoning, addressing quality degradation caused by fixed-length embeddings. Shen et al. (2025) employs self-distillation to align student and teacher hidden activations under CoT supervision, constraining latent reasoning paths.
In the multimodal domain, latent reasoning has also been introduced. Unlike traditional LLMs, VLMs emphasize how image features interact with the latent space. Some efforts Yang et al. (2025); Li et al. (2025a); Pham and Ngo (2025) have been made to integrate visual "thoughts" into the latent space for reasoning. However, these existing works focus solely on single-modal latent reasoning. Combining text and vision for multimodal latent reasoning in the latent space remains unexplored.
3 Method
@@ 3.1 Multimodal Latent Reasoning
Figure 2 provides an overview of our approach. In IVT-LR, the latent reasoning is conducted over both latent text and latent vision. Following Hao et al. (2024), the textual modality bypasses explicit token prediction: instead of using the embedding of the previous explicit text token, we represent the latent text with the hidden state $h_{t-1}^{hidden}$. Meanwhile, the latent vision is designed to model the dynamic focus on the visual features at each step. Specifically, we extend latent reasoning to visual modality by selecting the $k$ most relevant visual features from the image embedding set. Thus, an attention-based selection mechanism is designed to choose a fixed number of image embeddings from the full set $z_{1},z_{2},\dots,z_{J}]$. We utilize the sum of attention weights across all layers to identify the $k$ image embedding positions with the highest cumulative scores. The selected features are appended to the hidden states $h_{t-1}^{hidden}$, resulting in a multimodal latent representation $h_{t-1}^{\text{latent}},z_{t-1}^{\text{selected}}]$. The input to the model at step $t$ thus consists of all prior hidden states and their selected visual features, along with any preceding question embeddings, which can be written as $E_{t}=[e_{1},\dots,e_{N},h_{1}^{\text{latent}},z_{1}^{\text{selected}},\dots,h_{t-1}^{\text{latent}},z_{t-1}^{\text{selected}}]$. The model fuses these multimodal representations to obtain $e_{t}^{\text{fused}}=f(E_{t})$, which is projected through the output head to yield the next-token distribution $\mathcal{M}(x_{t+1}\mid E_{t})=\text{softmax}(W\cdot e_{t}^{\text{fused}})$. This design allows the model to perform step-wise multimodal latent reasoning without generating intermediate reasoning sequences.
@@
The objective of IVT-LR is to enable multimodal reasoning within the latent space. Inspired by  Deng et al. (2024), we adopt a multi-stage training strategy to progressively boost the model’s reasoning capability. In the preprocessing stage, each reasoning trajectory is segmented into up to $N$ steps, followed by the final answer. At stage 0, as shown in Figure 3, the model is trained with standard CoT supervision, where all reasoning steps are explicitly generated to strengthen symbolic reasoning ability. Afterwards, the latent reasoning steps are progressively introduced within the $N$ stages: at each stage, one additional explicit reasoning step is replaced by a latent reasoning step, denoted by the special token <latent>, beginning with the first step. In this way, the model learns to progressively substitute explicit reasoning using latent textual and visual representations while still being supervised on the final answer.
Training is optimized using negative log-likelihood (NLL) loss, with supervision applied only to reasoning steps and the final answer. Latent reasoning steps and question tokens are masked out. This design ensures that the supervision signal is placed only on the reasoning steps and the final answer, guiding the model to anticipate subsequent inference rather than redundantly imitating the replaced explicit step. By avoiding excessive alignment between latent representations and explicit rationales, the model learns to internalize reasoning trajectories in latent space with essential image features, while still being driven toward correct final predictions.
@@ 4 Experiments
Backbone Methods M3CoT ScienceQA
Qwen2-VL No-CoT 45.4 - - 64.4 - -
Multimodal CoT(Zhang et al., 2024) 42.5 106.3 3.10 58.3 83.9 2.44
CCoT(Mitra et al., 2024) 44.1 177.2 5.31 63.8 164.0 5.23
ICoT(Gao et al., 2025) 46.0 96.5 2.86 65.4 77.4 2.28
SCAFFOLD(Lei et al., 2025) 44.9 170.8 5.14 62.5 162.3 4.91
Chain-of-Focus(Zhang et al., 2025) 64.3 185.7 2.63 91.2 162.3 2.09
IVT-LR 71.8 10.0 0.65 94.6 11.0 0.67
Chameleon No-CoT 28.4 - - 48.5 - -
Multimodal CoT(Zhang et al., 2024) 30.6 110.5 3.62 50.7 98.7 3.33
CCoT(Mitra et al., 2024) 31.4 168.4 5.35 51.3 174.2 5.39
ICoT(Gao et al., 2025) 32.3 110.9 5.43 53.4 92.4 4.62
SCAFFOLD(Lei et al., 2025) 31.1 194.3 6.12 47.5 160.6 6.03
Chain-of-Focus(Zhang et al., 2025) 36.5 739.4 3.09 61.2 717.1 2.56
IVT-LR 41.8 10.0 1.13 64.0 11.0 1.14
@@
We evaluate our method on two widely used multimodal reasoning benchmarks: M3CoT (Chen et al., 2024) and ScienceQA (Lu et al., 2022). M3CoT is a large-scale benchmark focusing on multimodal chain-of-thought reasoning, where models must combine both visual and textual inputs to perform multi-step reasoning. ScienceQA is a diverse dataset covering natural science, language science, and social science, with many questions accompanied by diagrams or images. We evaluate using exact-match answer accuracy, along with the average number of autoregressive steps and the average response time per question. These metrics capture both correctness and reasoning efficiency.
@@
We compare IVT-LR against six representative methods, including text-only reasoning: CCoT (Mitra et al., 2024); vision-text involved reasoning: Chain-of-Focus (Zhang et al., 2025), SCAFFOLD (Lei et al., 2025), ICoT (Gao et al., 2025), Multimodal-CoT (Zhang et al., 2024); and No-CoT that directly predicts answers without generating intermediate steps.
For fair comparison, we evaluate IVT-LR and all baselines with Qwen2-VL-7B (Wang et al., 2024a) and Chameleon-7B (Team, 2024) backbones. In IVT-LR training, we use a stage number ($N$) of four (detailed discussion provided in Appendix A) , a batch size of four, and train with the Adam optimizer where the learning rate is set to $\mathbf{4\times 10^{-5}}$ and $\mathbf{\beta_{1}}$ is set to $0.9$. All experiments run on four NVIDIA A6000 GPUs (48GB VRAM each).

### v2

@@ Abstract
Multimodal reasoning aims to enhance the capabilities of MLLMs by incorporating intermediate reasoning steps before reaching the final answer. It has evolved from text-only reasoning to the integration of visual information, enabling the thought process to be conveyed through both images and text. Despite its effectiveness, current multimodal reasoning methods depend on explicit reasoning steps that require labor-intensive vision-text annotations and inherently introduce significant inference latency. To address these issues, we introduce multimodal latent reasoning with the advantages of multimodal representation, reduced annotation, and inference efficiency. To facilitate it, we propose Interleaved Vision-Text Latent Reasoning (IVT-LR), which injects both visual and textual information in the reasoning process within the latent space. Specifically, IVT-LR represents each reasoning step by combining two implicit parts: latent text (the hidden states from the previous step) and latent vision (a set of selected image embeddings). We further introduce a progressive multi-stage training strategy to enable MLLMs to perform the above multimodal latent reasoning steps. Experiments on M3CoT and ScienceQA demonstrate that our IVT-LR method achieves an average performance increase of 5.45% in accuracy, while simultaneously achieving a speed increase of over 5 times compared to existing approaches.
@@ 1 Introduction
Over the past few years, the capabilities of large language models (LLMs) have been further unlocked through advancements in reasoning. Researchers have sought to enhance the reasoning abilities of LLMs, initially through prompting techniques such as Chain-of-Thought prompting [wei2022chain], and more recently by developing large reasoning models using reinforcement learning, like GPT-4o [hurst2024gpt] and Deepseek-R1 [guo2025deepseek]. Building on the success of reasoning in LLMs, there has been growing interest in the research community to extend these reasoning capabilities to multimodal LLMs (MLLMs). This has brought the promising topic of multimodal reasoning, aiming to improve the performance of models on multimodal tasks, such as VQA, through reasoning.
@@
Current multimodal reasoning approaches can be broadly categorized into the following progressive steps: 1) Text-only reasoning. Early multimodal reasoning methods primarily focused on pure text-based reasoning, where MLLMs perform textual reasoning before generating the final answer. These approaches [hu2022promptcap, kamcot] could seamlessly apply LLMs methodologies to MLLMs. 2) Vision-text involved reasoning. Some studies [visualsketchpad, liu2025Visual, chern2025thinking] highlight that the intermediate reasoning steps also require the involvement of visual information. For instance, gao2024interleaved enhances reasoning by generating sequential steps that interleave visual information with textual rationales via selected image patches. In a related direction, zheng2025deepeyes trains models through end-to-end reinforcement learning to autonomously zoom in on image regions for fine-grained visual inspection during the reasoning process. Alternatively, li2025imaginereasoningspacemultimodal enables MLLMs to actively “think visually” by generating explicit image visualizations of their reasoning traces, thereby significantly enhancing performance on complex spatial reasoning tasks.
Recently, latent reasoning has emerged as a new paradigm in LLMs, which eliminates the need for explicit and lengthy textual reasoning by leveraging implicit latent vectors [hao2024training]. Inspired by this, we believe that latent reasoning holds even greater potential for facilitating vision-text interleaved intermediate reasoning steps due to the following reasons: 1) Multimodal representation potential. Latent reasoning enables the reasoning process to occur entirely within a hidden space, offering a greater capacity to represent rich, multimodal information during reasoning. 2) Reduced Annotation. Introducing latent reasoning will lessen reliance on heavily annotated vision-text interleaved reasoning data, as reasoning steps no longer need to be fully observable or linguistically aligned. 3) Inference efficiency. By avoiding long chains of explicit multimodal representation in the reasoning step, it will significantly improve efficiency.
@@
  • We introduce IVT-LR, the first framework to achieve fully unified multimodal latent reasoning. Unlike prior methods, our approach enables both textual and visual information to be reasoned with in the latent space, eliminating the need for intermediate explicit text or image generation.
@@ 2.1 Multimodal Reasoning
Text-only reasoning.
Early works attempt to convert visual information into text before reasoning, using tools or visual experts to generate textual representations to guide LLMs. hu2022promptcap first introduced the concept of captions, extracting visual content as textual captions and concatenating them to the input to enhance reasoning. Inspired by this, subsequent works pursued a fine-grained understanding of images to improve textual expressiveness. Zheng_NeurIPS2023 generates a rationale that incorporates image information from visual-text inputs, which is then used for reasoning. Other works [MitraCCoT, kamcot] leverage graph structures to identify entities in images and construct relationships among them, enhancing reasoning based on these inter-entity connections.
Vision-text involved reasoning.
This line of work emphasizes using images together with text during the rationale generation and reasoning process. Building on the reasoning paradigm of large language models, zhang2023multimodal first proposed decoupling rationale generation from answer generation in the Vision-Text Reasoning field. Subsequently, Shao2024VisualCA annotates key regions of the original image in intermediate steps, training models to focus on image regions relevant to the answer. While some works [gao2024interleaved, zhang2025chain] further extract key image regions progressively during reasoning, combining visual information with textual reasoning to generate the final answer. Moreover, new methods [visualsketchpad, liu2025Visual] emulate human thought by sketching images during reasoning, focusing on core concepts, structures, and relationships while ignoring redundant details. Other works [li2025imaginereasoningspacemultimodal, chern2025thinking] generate new images during reasoning, combining them with text to improve reasoning in complex scenarios. To completely decouple reasoning from language and amplify the role of images, xu2025visualplanningletsthink proposes reasoning solely with newly generated images, achieving substantial improvements in visual navigation tasks.
2.2 Latent Reasoning
Latent reasoning refers to internal, non-linguistic thinking performed in a hidden latent space before generating the final answer. Early methods used special tokens to guide latent reasoning. goyal2023think introduces learnable <pause> tokens, giving the model opportunities to internally update information before generating an answer, while wang2023guiding uses <plan> tokens to guide reasoning.
Later, some works exploit the model’s continuous hidden states to replace explicit reasoning steps. hao2024training pioneers continuous latent space reasoning by feeding the last hidden states as input embeddings for the next step without generating intermediate tokens, significantly reducing reasoning tokens and improving efficiency. Inspired by this, subsequent methods improve the quality of intermediate representations. cheng2024compressed uses variable-length contemplation tokens for latent reasoning, addressing quality degradation caused by fixed-length embeddings. shen2025codi employs self-distillation to align student and teacher hidden activations under CoT supervision, constraining latent reasoning paths.
In the multimodal domain, latent reasoning has also been introduced. Unlike traditional LLMs, VLMs emphasize how image features interact with the latent space. Some efforts [yang2025machine, li2025latent, pham2025multimodal] have been made to integrate visual "thoughts" into the latent space for reasoning. However, these existing works focus solely on single-modal latent reasoning. Combining text and vision for multimodal latent reasoning in the latent space remains unexplored.
3 Method
@@ 3.1 Multimodal Latent Reasoning
Figure 2 provides an overview of our approach. In IVT-LR, the latent reasoning is conducted over both latent text and latent vision. Following [hao2024training], the textual modality bypasses explicit token prediction: instead of using the embedding of the previous explicit text token, we represent the latent text with the hidden state $h_{t-1}^{hidden}$, which effectively encodes the necessary reasoning logic and preserves richer intermediate information in a continuous latent space. Meanwhile, the latent vision is designed to model the dynamic focus on the visual features at each step. Specifically, we extend latent reasoning to visual modality by selecting the $k$ most relevant visual features from the image embedding set. Thus, an attention-based selection mechanism is designed to choose a fixed number of image embeddings from the full set $z_{1},z_{2},\dots,z_{J}]$. We utilize the sum of attention weights across all layers to identify the $k$ image embedding positions with the highest cumulative scores. The selected features are appended to the hidden states $h_{t-1}^{hidden}$, resulting in a multimodal latent representation $h_{t-1}^{\text{latent}},z_{t-1}^{\text{selected}}]$. The input to the model at step $t$ thus consists of all prior hidden states and their selected visual features, along with any preceding question embeddings, which can be written as $E_{t}=[e_{1},\dots,e_{N},h_{1}^{\text{latent}},z_{1}^{\text{selected}},\dots,h_{t-1}^{\text{latent}},z_{t-1}^{\text{selected}}]$. The model fuses these multimodal representations to obtain $e_{t}^{\text{fused}}=f(E_{t})$, which is projected through the output head to yield the next-token distribution $\mathcal{M}(x_{t+1}\mid E_{t})=\text{softmax}(W\cdot e_{t}^{\text{fused}})$. This design allows the model to perform step-wise multimodal latent reasoning without generating intermediate reasoning sequences.
@@
The objective of IVT-LR is to enable multimodal reasoning within the latent space. Inspired by deng2024implicitcot, we adopt a multi-stage training strategy to progressively boost the model’s reasoning capability. In the preprocessing stage, each reasoning trajectory is segmented into up to $N$ steps, followed by the final answer. At stage 0, as shown in Figure 3, the model is trained with standard CoT supervision, where all reasoning steps are explicitly generated to strengthen symbolic reasoning ability. Afterwards, the latent reasoning steps are progressively introduced within the $N$ stages: at each stage, one additional explicit reasoning step is replaced by a latent reasoning step, denoted by the special token <latent>, beginning with the first step. In this way, the model learns to progressively substitute explicit reasoning using latent textual and visual representations while still being supervised on the final answer.
Training is optimized using negative log-likelihood (NLL) loss, with supervision applied only to reasoning steps and the final answer. Latent reasoning steps and question tokens are masked out. This design ensures that the supervision signal is placed only on the reasoning steps and the final answer, distinguishing our approach from standard knowledge distillation. Unlike distillation, which enforces alignment between student hidden states and a teacher’s explicit reasoning trajectory, our method avoids imposing a strict linear path. By avoiding excessive alignment between latent representations and explicit rationales, the model learns to internalize reasoning trajectories in latent space with essential image features, while still being driven toward correct final predictions. Compared to single-step fine-tuning, the proposed multi-stage training introduces additional stages but reuses the same backbone and training data, resulting in a moderate increase in training cost that remains practical and acceptable.
Backbone Methods M3CoT ScienceQA
Qwen2-VL No-CoT 45.4 - - 64.4 - -
Multimodal CoT[zhang2023multimodal] 42.5 106.3 3.10 58.3 83.9 2.44
CCoT[MitraCCoT] 44.1 177.2 5.31 63.8 164.0 5.23
ICoT[gao2024interleaved] 46.0 96.5 2.86 65.4 77.4 2.28
SCAFFOLD[lei2024scaffolding] 44.9 170.8 5.14 62.5 162.3 4.91
Chain-of-Focus[zhang2025chain] 64.3 185.7 2.63 91.2 162.3 2.09
IVT-LR 71.8 10.0 0.65 94.6 11.0 0.67
Chameleon No-CoT 28.4 - - 48.5 - -
Multimodal CoT[zhang2023multimodal] 30.6 110.5 3.62 50.7 98.7 3.33
CCoT[MitraCCoT] 31.4 168.4 5.35 51.3 174.2 5.39
ICoT[gao2024interleaved] 32.3 110.9 5.43 53.4 92.4 4.62
SCAFFOLD[lei2024scaffolding] 31.1 194.3 6.12 47.5 160.6 6.03
Chain-of-Focus[zhang2025chain] 36.5 739.4 3.09 61.2 717.1 2.56
IVT-LR 41.8 10.0 1.13 64.0 11.0 1.14
@@ 4 Experiments
We evaluate our method on two widely used multimodal reasoning benchmarks: M3CoT [chen-etal-2024-m3cot] and ScienceQA [lu2022learn]. M3CoT is a large-scale benchmark focusing on multimodal chain-of-thought reasoning, where models must combine both visual and textual inputs to perform multi-step reasoning. ScienceQA is a diverse dataset covering natural science, language science, and social science, with many questions accompanied by diagrams or images. We evaluate using exact-match answer accuracy, along with the average number of autoregressive steps and the average response time per question. These metrics capture both correctness and reasoning efficiency.
@@
We compare IVT-LR against six representative methods, including text-only reasoning: CCoT [MitraCCoT]; vision-text involved reasoning: Chain-of-Focus [zhang2025chain], SCAFFOLD [lei2024scaffolding], ICoT [gao2024interleaved], Multimodal-CoT [zhang2023multimodal]; and No-CoT that directly predicts answers without generating intermediate steps.
For fair comparison, we evaluate IVT-LR and all baselines with Qwen2-VL-7B [wang2024qwen2] and Chameleon-7B [team2024chameleon] backbones. In IVT-LR training, we use a stage number ($N$) of four (detailed discussion provided in Appendix A) , a batch size of four, and train with the Adam optimizer where the learning rate is set to $\mathbf{4\times 10^{-5}}$ and $\boldsymbol{\beta}_{1}$ is set to $0.9$. All experiments run on four NVIDIA A6000 GPUs (48GB VRAM each).